# TikTok’s recommendations skewed towards Republican content during the 2024 U.S. presidential race

### v1

@@ Introduction
Yet, despite its growing influence and the ongoing discussion surrounding its ban, TikTok remains relatively unexplored in the literature on social media recommendation algorithms. There is extensive prior work auditing social media recommendation algorithms and user behavior on platforms such as Facebook [18, 19, 20, 21, 22, 23, 24], Instagram [25, 26, 27, 28], YouTube [29, 30, 31, 32, 33], X (formerly Twitter) [34, 35, 36, 37, 38, 39], and Reddit [40, 41, 42], but little on TikTok, potentially due to its relative novelty as a social media platform and tight restrictions on its official API. Nevertheless, some important studies have examined political content on TikTok through different lenses, such as investigating political user demographics [43], the characteristics of content that political parties publish on the platform [44, 45, 46, 47], polarization and toxicity in TikTok political discourse [48], and the intersection of humor and political engagement [49, 50].
Nonetheless, this literature leaves two important gaps. We know little about the number and partisan distribution of political content creators on TikTok, or their engagement and audiences; we know even less about TikTok’s recommendation algorithm. We begin to fill these gaps here.
@@
We conduct a series of 323 audit experiments on the platform using sock puppets (i.e., bots), mimicking user activity. Each experiment employs a bot to watch a researcher-controlled sequence of videos seeded with political content, which we call the “conditioning stage.” This treatment conditioning teaches TikTok’s recommendation algorithm our bot’s political preferences. Following the conditioning stage, our bot watches videos on TikTok’s recommendation page (the “For You” page), which we call the “recommendation stage.” In this stage, each bot repeats the process of watching 10 videos, followed by a one-hour pause in which the bot does not engage with TikTok (to circumvent TikTok’s bot-detection algorithm). The content of these videos constitutes our outcome data: we infer and analyze the partisan, topical, and misinformation content of these videos, including partisan alignment between the conditioning videos and the recommended ones. In total, each experimental run was conducted over the duration of one week, with one day dedicated to the conditioning stage, and six days dedicated to the recommendation stage (see Supplementary Figure S2 for an illustration of a single experimental run).
@@
A bot’s geography is the location where the bot is virtually present. Each bot is connected to a VPN server in New York, Texas, or Georgia. These were selected due to their projection as a likely Democratic, Republican, or swing state in the 2024 U.S. elections, respectively. In addition to VPN tunneling, each bot used a GPS mocking application to artificially set its location to its given VPN state (for more details, see the Pre- and Post-Experiment Protocols section in the Supplementary material).
@@
We exclude a number of experimental runs for technical reasons: some bots were shut down by TikTok’s bot detection algorithm, while others failed due to internet outages. To account for this and ensure that we compare bots with the same amount of conditioning and the same number of recommended videos, we conduct a pair-matching procedure to match bots of opposite conditioning in each state during a given week. For each pair of bots, we only consider the first N recommendations, where N is the lesser of the total videos watched by either bot. We also exclude pairs of bots that watched fewer than 150 videos each during the recommendation stage to allow for a sufficient number of political videos to be recommended, bearing in mind that not all recommendations are political in nature. Of the 567 experimental runs conducted, 323 met our inclusion criteria. Supplementary Table S11 lists the number of successful experimental runs per condition over time.
Through 323 experiment runs over 27 weeks, our bots watched 114,522 total (3,181 unique) videos in the conditioning stage and an additional 284,093 (176,252 unique) videos in the recommendation stage. Figure 1 illustrates our experimental design and data collection procedure; more details are available in the Materials and Methods section of the Supplementary Materials. The study was preregistered shortly before the end of the data collection period, and before the data was analyzed (asPredicted link: https://aspredicted.org/mjj2-vxrb.pdf).
@@
We first download the English transcripts of each video from the subtitle URLs provided in their metadata, which were available in 40,264 videos, constituting 22.8% of the unique videos and 26.2% of all videos recommended to our bots. To categorize a given video’s partisan content, we use a human-validated pipeline utilizing three large language models (LLMs)—GPT-4o [53], Gemini-Pro [54], and GPT-4 [55]—to answer the following questions about a given video: (Q1) Is the video political in nature?, (Q2) Is the video concerned with the 2024 U.S. elections or major U.S. political figures?, and (Q3) What is the ideological stance of the video in question (Pro Democratic, Anti Democratic, Pro Republican, Anti Republican, or Neutral)? For each video, we prompt each LLM to answer Q1, and if the answer is Yes, we ask Q2 and Q3. For each question, our outcome measure is the majority vote of the three LLMs’ answers. A majority is guaranteed for Q1 and Q2 as they have binary outcomes. As for Q3, despite having five outcome categories, 89.4% of videos reached a majority label; our analysis focuses on videos with an LLM majority vote. In the Data Representativeness section of the Supplementary materials, we show that the partisan distributions of videos with transcripts does not differ significantly from a large sample of videos without transcripts, indicating that the videos analyzed in this study are representative of the entire set of recommendations made to the bots.
The above classification pipeline improves upon prior work characterizing video-level political content on TikTok. Tjaden et al., who study TikTok political content in the context of German elections [56], associate a video with a party if it includes a party-affiliated hashtag in its description. This approach is not applicable in the U.S., where videos which include hashtags containing “trump” are often published by Democratic-aligned creators. Notably, “#donaldtrump” and “#trump” are among the six most common hashtags used by Democratic-aligned channels (Supplementary Table S19). We provide more details on our measurement, validation, and prompting in the Transcript Classification section of the Supplementary Materials, and we present descriptive statistics about our complete dataset in Supplementary Table S12.
@@
What is the supply and partisan distribution of political content creators on TikTok? To categorize channels as Democratic-aligned or Republican-aligned, we calculate the proportion of each creator’s videos labeled as Pro-Democratic or Anti-Republican versus Pro-Republican or Anti-Democratic, supplementing our dataset with up to 30 additional pre-election videos from the TikAPI [57] for channels with fewer than 10 labeled videos in our sample. We label a channel as Democratic-aligned if at least 75% of its videos are either Pro-Democratic or Anti-Republican, and Republican-aligned if at least 75% of its videos are Pro-Republican or Anti-Democrat. This process yielded 56 Democratic-aligned channels and 75 Republican-aligned channels, which we manually validated following best practices on channel-level classification tasks  [30, 32]. Supplementary Table S8 summarizes the average proportion of party-aligned videos across these channels.
@@
We measure the “ideological content” of a given bot’s recommendations as the proportion of political recommended videos that are Republican-aligned (Pro Republican or Anti Democratic) minus the proportion of recommendations that are Democratic-aligned (Pro Democratic or Anti Republican). This allows us to score each bot’s video recommendations from -1 (all videos are Democratic-aligned) to +1 (all videos are Republican-aligned). We plot the bots’ recommendations across experimental conditions and over time in Figure 2A, B, C for the states of New York, Texas, and Georgia, respectively. Each point represents the mean ideological content of the recommendations received by the bots in a given week and experimental condition. Here, bars indicate 95% confidence intervals, while color indicates the conditioned partisanship of the bots. Missing points represent weeks without any successful runs for bots located in a given state (see Supplementary Table S11 for the number of successful runs per condition over time). We also plot OLS regression lines of ideological content by experimental condition over time, showing a clear increase in polarization in all states: TikTok served partisans more co-partisan content over time. However, while both Democratic and Republican-conditioned bots received more ideologically aligned content over time, the magnitude is significantly greater for Republicans.
@@
We confirm that the TikTok algorithm exhibits an ideological skew towards Republican-aligned content in Figure 2D, E, and F. In these figures, the Y axis shows the overall partisan content of all TikTok’s recommended videos regardless of partisan alignment, where negative values indicate a Democratic bias and positive values indicate a Republican bias. On average, Republican bots saw approximately 11.8% more co-partisan recommendations than their Democratic bot counterparts while simultaneously seeing approximately 7.5% less cross-partisan recommendations. Supplementary Table S13 details the mean monthly ideological content of Republican and Democratic bots in each state, as well as Neutral bots in Georgia. Supplementary Table S14 details results from independent t-tests comparing the partisan content of bots of each type of conditioning, revealing that Republican bots received significantly more co-partisan content across all three states (NY: $t=5.82$, $p<0.001$; Texas: $t=9.21$, $p<0.001$; Georgia: $t=5.83$, $p<0.001$), and both Republican and Democratic bots received more co-partisan content than Neutral bots (Rep vs. Neutral: $t=13.22$, $p<0.001$; Dem vs. Neutral: $t=4.70$, $p<0.001$). Supplementary  Figures S3, S4, S5 and S6 illustrate rolling averages of the amount of political content, election-related content, ideological matches, and ideological mismatches viewed by different bots during a given experimental run. As shown in these figures, for both Democratic- and Republican-conditioned bots, the conditioning stage results in an immediate spike in political content recommended to bots, decaying as bots are recommended more videos but never approaching the control group.
@@
To verify that our results are not due to differences in the engagement metrics of Republican and Democratic videos or channels, we consider counterfactual scenarios where video recommendations are functions of these engagement metrics. We showed above that the TikTok algorithm recommends more Republican-aligned content than Democratic-aligned content, but this may not be surprising if there is more Republican content overall on TikTok, or if that content is more popular. Our robustness tests answer the question: how big of an ideological skew should we expect under different scenarios, and how does the observed skew compare?
@@
To confirm that the skew towards Republican-aligned content exists even after accounting for potential differences in video engagement metrics, we take a weighted-random sample of N videos (N being the number of videos watched by pairs of bots in a given week and experimental condition) with weights proportional to that video’s engagement, and calculate the proportion of Republican- and Democratic-aligned videos in that sample. We then compare these proportions to the observed proportion of recommended Republican and Democratic-aligned videos, and show that our bots received more such Republican-aligned videos than we would expect if recommendations were only a function of video engagement. To this end, we consider 39 scenarios where recommendation rates are a function of TikTok engagement metrics: video-level metrics such as a video’s likes, likes per recommendation, comments, comments per recommendation, plays, and share count, as well as channel-level metrics such as verification status, follower count, and cumulative like and video counts. We aggregate these metrics using principal component analysis (PCA) and extract the video loadings to produce a combined video engagement score, a combined channel engagement score, and a combined overall engagement score. For each of the 13 models described (six video-level models, four channel-level models, and three combined models), we also compute two alternative versions, penalizing old videos linearly and exponentially, to up-weight more recent content.
@@
We show the results from this counterfactual analysis in Figure 2G. The leftmost point represents the observed ideological skew (the proportion of Republican-aligned videos minus that of Democratic-aligned ones recommended to the bots). Each of the remaining points presents the expected ideological skew as computed by the 39 counterfactuals, shown along the x-axis; plot point markers represent the alternative methods for up-weighting more recent content. Across all 39 counterfactuals, the observed bias toward Republican-aligned content substantially exceeded the counterfactual proportion based on engagement. Supplementary Table S15 lists the expected ideological skews as computed by each model, as well as independent t-tests comparing these values to that observed by the bots; see the Video Recommendation Counterfactual Models section of the Supplementary Materials for further details.
@@
We formalize this analysis in a logistic regression framework (Figure 2H). We conduct a regression at the bot-video recommendation level where the outcome variable indicates whether a video is an ideological mismatch as a function of that bot’s conditioned partisanship, controlling for that bot’s state, the week in which the experiment was conducted, as well as the engagement metrics of the videos and channels watched during the recommendation stage, in addition to those watched in the conditioning stage. We isolate the independent variables with a Variance Inflation Factor (VIF) value less than five (Supplementary Table S21 lists the VIF scores for each of the independent variables). Considering that all conditioning video- and channel-level engagement metrics had a VIF score greater than five, we compute a combined engagement score for those videos using PCA. In Figure 2H, the top-most point shows the logistic regression coefficient on bot conditioning party. As can be seen, compared to Republican-conditioned bots, Democratic-conditioned bots are 2.8 times as likely to receive ideologically mismatched videos. We show the full regression table for this model in Supplementary Table S23. Additionally, Supplementary Tables S26, S27, S28 and S29 include descriptive statistics of the engagement metrics of videos and channels seen by bots during the conditioning and recommendation stages.
@@
Indeed, of the top Republican and Democratic channels, the highest four mismatch proportions were of Republican channels (daterightstuff, foxnews, realdonaldtrump, and teamtrump). Overall, Republican channels had a significantly higher mismatch proportion than Democratic channels (Chi-squared test; $\chi^{2}=36.5$, $p<0.001$). Figure 3B shows that the official channels of the Republican candidates, Donald Trump and JD Vance, had a significantly higher mismatch proportion than those of Democratic candidates Kamala Harris and Tim Walz (Chi-squared test; $\chi^{2}=28.1$, $p<0.001$). This partisan discrepancy in the mismatch proportion observed between these channels is most strongly associated with the bot’s partisan conditioning (Figure 3C). For this analysis, we again considered variables with a VIF value less than 5, as listed in Supplementary Table S22. Democratic-conditioned bots were roughly 1.75 times more likely to view videos made by Republican channels than Republican bots were to view videos by Democratic channels. See Supplementary Table S24 for the logistic regression table containing these results.

### v2

@@ Introduction
Yet, despite its growing influence and the ongoing discussion surrounding its ban, TikTok remains relatively unexplored in the literature on social media recommendation algorithms. There is extensive prior work auditing social media recommendation algorithms and user behavior on platforms such as Facebook [18, 19, 20, 21, 22, 23, 24], Instagram [25, 26, 27, 28], YouTube [29, 30, 31, 32, 33], X (formerly Twitter) [34, 35, 36, 37, 38, 39], and Reddit [40, 41, 42], but little on TikTok, potentially due to its relative novelty as a social media platform and tight restrictions on its official API. Nevertheless, some important studies have examined political content on TikTok through different lenses, such as investigating political user demographics [43], the characteristics of content that political parties publish on the platform [44, 45, 46, 47], polarization and toxicity in TikTok political discourse [48], and the intersection of humor and political engagement [49, 50]. TikTok specifically is uniquely positioned as a platform of interest due to the manner in which users engage with the service. On YouTube, for instance, while users are offered recommendations (which is impacted by both the user’s own preferences, as well as co-viewership networks), the choice of the next video a user watches largely lies with the user (excluding YouTube shorts, or allowing videos to auto-play), which we term as “direct” agency. On TikTok, while users can potentially alter their feed by engaging with videos in various ways (e.g. liking, commenting on, or sharing a video), users ultimately only have “indirect” agency with regards to what video they watch next on their feed. As such, TikTok’s recommendation algorithm plays a far more central role in what users consume relative to other social media services.
Taken together, we believe the literature leaves two important gaps. We know little about the number and partisan distribution of political content creators on TikTok, or their engagement and audiences; we know even less about TikTok’s recommendation algorithm. We begin to fill these gaps here.
@@
We conduct a series of 323 audit experiments on the platform using sock puppets (i.e., bots), mimicking user activity. Each experiment employs a bot to watch a researcher-controlled sequence of videos seeded with political content, which we call the “conditioning stage.” This treatment conditioning teaches TikTok’s recommendation algorithm our bot’s political preferences. Following the conditioning stage, our bot watches videos on TikTok’s recommendation page (the “For You” page), which we call the “recommendation stage.” In this stage, each bot repeats the process of watching 10 videos, followed by a one-hour pause in which the bot does not engage with TikTok (to circumvent TikTok’s bot-detection algorithm). The content of these videos constitutes our outcome data: we infer and analyze the partisan, topical, and misinformation content of these videos, including partisan alignment between the conditioning videos and the recommended ones. In total, each experimental run was conducted over the duration of one week, with one day dedicated to the conditioning stage, and six days dedicated to the recommendation stage (see Supplementary Figure 1 for an illustration of a single experimental run).
@@
A bot’s geography is the location where the bot is virtually present. Each bot is connected to a VPN server in New York, Texas, or Georgia. These were selected due to their projection as a likely Democratic, Republican, or swing state in the 2024 U.S. elections, respectively. In addition to VPN tunneling, each bot used a GPS mocking application to artificially set its location to its given VPN state (for more details, see the Pre- and Post-Experiment Protocols in Methods).
@@
We exclude a number of experimental runs for technical reasons: some bots were shut down by TikTok’s bot detection algorithm, while others failed due to internet outages. To account for this and ensure that we compare bots with the same amount of conditioning and the same number of recommended videos, we conduct a pair-matching procedure to match bots of opposite conditioning in each state during a given week. For each pair of bots, we only consider the first N recommendations, where N is the lesser of the total videos watched by either bot. We also exclude pairs of bots that watched fewer than 150 videos each during the recommendation stage to allow for a sufficient number of political videos to be recommended, bearing in mind that not all recommendations are political in nature. Of the 567 experimental runs conducted, 323 met our inclusion criteria. Supplementary Table 1 lists the number of successful experimental runs per condition over time.
Through 323 experiment runs over 27 weeks, our bots watched 114,522 total (3,181 unique) videos in the conditioning stage and an additional 284,093 (176,252 unique) videos in the recommendation stage. Figure 1 illustrates our experimental design and data collection procedure; more details are available in the Experimental Setup section in Methods. The study was preregistered shortly before the end of the data collection period, and before the data was analyzed (asPredicted link: https://aspredicted.org/mjj2-vxrb.pdf).
@@
We first download the English transcripts of each video from the subtitle URLs provided in their metadata, which were available in 40,264 videos, constituting 22.8% of the unique videos and 26.2% of all videos recommended to our bots. To categorize a given video’s partisan content, we use a human-validated pipeline utilizing three large language models (LLMs)—GPT-4o [53], Gemini-Pro [54], and GPT-4 [55]—to answer the following questions about a given video: (Q1) Is the video political in nature?, (Q2) Is the video concerned with the 2024 U.S. elections or major U.S. political figures?, and (Q3) What is the ideological stance of the video in question (Pro Democratic, Anti Democratic, Pro Republican, Anti Republican, or Neutral)? For each video, we prompt each LLM to answer Q1, and if the answer is Yes, we ask Q2 and Q3. For each question, our outcome measure is the majority vote of the three LLMs’ answers. A majority is guaranteed for Q1 and Q2 as they have binary outcomes. As for Q3, despite having five outcome categories, 89.4% of videos reached a majority label; our analysis focuses on videos with an LLM majority vote. In the Data Representativeness in Methods, we show that the partisan distributions of videos with transcripts does not differ significantly from a large sample of videos without transcripts, indicating that the videos analyzed in this study are representative of the entire set of recommendations made to the bots.
The above classification pipeline improves upon prior work characterizing video-level political content on TikTok. Tjaden et al., who study TikTok political content in the context of German elections [56], associate a video with a party if it includes a party-affiliated hashtag in its description. This approach is not applicable in the U.S., where videos which include hashtags containing “trump” are often published by Democratic-aligned creators. Notably, “#donaldtrump” and “#trump” are among the six most common hashtags used by Democratic-aligned channels (Supplementary Table 9). We provide more details on our measurement, validation, and prompting in the Ideological Stance Classification section in Methods, and we present descriptive statistics about our complete dataset in Supplementary Table 2.
@@
What is the supply and partisan distribution of political content creators on TikTok? To categorize channels as Democratic-aligned or Republican-aligned, we calculate the proportion of each creator’s videos labeled as Pro-Democratic or Anti-Republican versus Pro-Republican or Anti-Democratic, supplementing our dataset with up to 30 additional pre-election videos from the TikAPI [57] for channels with fewer than 10 labeled videos in our sample. We label a channel as Democratic-aligned if at least 75% of its videos are either Pro-Democratic or Anti-Republican, and Republican-aligned if at least 75% of its videos are Pro-Republican or Anti-Democrat. This process yielded 56 Democratic-aligned channels and 75 Republican-aligned channels, which we manually validated following best practices on channel-level classification tasks  [30, 32]. Table 9 in the Methods section summarizes the average proportion of party-aligned videos across these channels.
@@
We measure the “ideological content” of a given bot’s recommendations as the proportion of political recommended videos that are Republican-aligned (Pro Republican or Anti Democratic) minus the proportion of recommendations that are Democratic-aligned (Pro Democratic or Anti Republican). This allows us to score each bot’s video recommendations from -1 (all videos are Democratic-aligned) to +1 (all videos are Republican-aligned). We plot the bots’ recommendations across experimental conditions and over time in Figure 2A, B, C for the states of New York, Texas, and Georgia, respectively. Each point represents the mean ideological content of the recommendations received by the bots in a given week and experimental condition. Here, bars indicate 95% confidence intervals, while color indicates the conditioned partisanship of the bots. Missing points represent weeks without any successful runs for bots located in a given state (see Supplementary Table 1 for the number of successful runs per condition over time). We also plot OLS regression lines of ideological content by experimental condition over time, showing a clear increase in polarization in all states: TikTok served partisans more co-partisan content over time. However, while both Democratic and Republican-conditioned bots received more ideologically aligned content over time, the magnitude is significantly greater for Republicans.
@@
We confirm that the TikTok algorithm exhibits an ideological skew towards Republican-aligned content in Figure 2D, E, and F. In these figures, the Y axis shows the overall partisan content of all TikTok’s recommended videos regardless of partisan alignment, where negative values indicate a Democratic bias and positive values indicate a Republican bias. On average, Republican bots saw approximately 11.8% more co-partisan recommendations than their Democratic bot counterparts while simultaneously seeing approximately 7.5% less cross-partisan recommendations. Supplementary Table 3 details the mean monthly ideological content of Republican and Democratic bots in each state, as well as Neutral bots in Georgia. Supplementary Table 4 details results from independent t-tests comparing the partisan content of bots of each type of conditioning, revealing that Republican bots received significantly more co-partisan content across all three states (NY: $t=5.82$, $p<0.001$; Texas: $t=9.21$, $p<0.001$; Georgia: $t=5.83$, $p<0.001$), and both Republican and Democratic bots received more co-partisan content than Neutral bots (Rep vs. Neutral: $t=13.22$, $p<0.001$; Dem vs. Neutral: $t=4.70$, $p<0.001$). Supplementary Figures 2-5 illustrate rolling averages of the amount of political content, election-related content, ideological matches, and ideological mismatches viewed by different bots during a given experimental run. As shown in these figures, for both Democratic- and Republican-conditioned bots, the conditioning stage results in an immediate spike in political content recommended to bots, decaying as bots are recommended more videos but never approaching the control group.
@@
Figure 2 establishes that TikTok shows more Republican-aligned content than Democratic-aligned content, and that Democrats receive more Republican-aligned content than Republicans receive Democratic-aligned content. However, if there are more Republican users or content creators on TikTok, if Republican-aligned content is more engaging, or if Democrats are more interested in Republican-aligned content than Republicans are in Democratic-aligned videos, then this ideological skew is exactly what we might expect TikTok’s algorithm to produce. In this section we show that these mechanisms cannot fully explain the observed difference in ideological content. First, we use a weighted random sampling approach to show that observed engagement metrics like channel size, video likes, and commenting, or linear combinations of the above, would not produce the observed ideological skew. Second, we show that the number of Democratic-aligned comments on a Republican-aligned video does not predict whether that video is shown more to Democrats. Neither of these two explanations, however, preclude the possibility that TikTok has developed a latent engagement metric that would explain the observed ideological gap; our third robustness test develops a sensitivity analysis to show that for any hypothetical hidden engagement metric, the difference in that metric between Republican and Democratic videos would have to be 98 times bigger than the gap in likes to explain the ideological content gap we observe on TikTok in our experiments.
Sampling by Engagement Metrics
@@
To verify that our results are not due to differences in the engagement metrics of Republican and Democratic videos or channels, we consider counterfactual scenarios where video recommendations are functions of these engagement metrics. We showed above that the TikTok algorithm recommends more Republican-aligned content than Democratic-aligned content, but this may not be surprising if there is more Republican content overall on TikTok, or if that content is more popular. In other words, Figure 2 assumes a counterfactual where Republican-aligned and Democratic-aligned videos appear in equal proportion; we consider alternative counterfactuals below. Our first robustness test answers the question: how big of an ideological skew in content should we expect under different counterfactual scenarios, and how does the observed skew compare?
@@
To confirm that the skew towards Republican-aligned content exists even after accounting for potential differences in video engagement metrics that might produce baseline differences, we take a weighted-random sample of N videos (N being the number of videos watched by pairs of bots in a given week and experimental condition) with weights proportional to that video’s engagement, and calculate the proportion of Republican- and Democratic-aligned videos in that sample. We then compare these proportions to the observed proportion of recommended Republican and Democratic-aligned videos, and show that our bots received more such Republican-aligned videos than we would expect if recommendations were only a function of video engagement. To this end, we consider 48 scenarios where recommendation rates are a function of TikTok engagement metrics: video-level metrics such as a video’s likes, likes per recommendation, comments, comments per recommendation, plays, and share count, as well as channel-level metrics such as verification status, follower count, and cumulative like and video counts. We then aggregate the above metrics using principal component analysis (PCA) and extract the video loadings to produce a combined video engagement score, a combined channel engagement score, and a combined overall engagement score. For each of the 16 models described (nine video-level models, four channel-level models, and three combined models), we also compute two alternative versions, penalizing old videos linearly and exponentially, to up-weight more recent content.
@@
We show the results from this counterfactual analysis in Figure 2G. The leftmost point represents the observed ideological skew (the proportion of Republican-aligned videos minus that of Democratic-aligned ones recommended to the bots). Each of the remaining points presents the expected ideological skew as computed by the 48 counterfactuals, shown along the x-axis; plot point markers represent the alternative methods for up-weighting more recent content. If the observed engagement metrics fully explain the partisan gap, we would expect the simulated gaps when we sample proportionate to those metrics to be close to the observed gap of approximately 0.2. We do not observe that. Instead, across all 48 counterfactuals, the observed bias toward Republican-aligned content substantially exceeded the counterfactual proportion based on engagement, with the majority of engagement metrics yielding a Democratic bias rather than a Republican one. Supplementary Table 5 lists the expected ideological skews as computed by each model, as well as independent t-tests comparing these values to that observed by the bots; see the  Video Recommendation Counterfactual Models section in Methods for further details.
Asymmetric Homophily
In addition to these public engagement metrics, we also consider bias which may stem from co-viewership networks [hosseinmardi2024causally]. Republican-aligned videos may garner more engagement from Democrats than vice-versa, which would explain why Republican-aligned videos are recommended to Democrats more frequently; we refer to this possibility as asymmetric homophily. A key implication of asymmetric homophily is that among Republican-aligned videos, those with more Democratic-aligned comments will be shown to Democrats more than those with fewer Democratic-aligned comments. We show that is not the case, suggesting that asymmetric homophily cannot explain the ideological skew we observe.
To test this, we collect a random sample of 200 comments from all political videos examined in our study and classify the content of each comment in the context of that video. From this we compute the proportion of comments ideologically aligned with the video (see the Comment Classification section in Methods).
We find meaningful differences between Republican-aligned and Democratic-aligned videos in the composition of comments, but not in the expected direction. We find that Democratic-aligned videos have a somewhat higher proportion of Republican-aligned comments than vice versa (13.4% versus 10.3%, $\chi^{2}=0.299$, $p=0.59$), as well as a higher proportion of copartisan comments than Republican-aligned videos (64.7% versus 52.9%, $\chi^{2}=0.544$, $p=0.46$). However, Republican-aligned videos have a much higher proportion of “neutral” comments that are not explicitly partisan (45.6% versus 27%, $\chi^{2}=2.07$, $p=0.15$). We also include this metric in our weighted sampling robustness check (Figure 2G), finding similar results to our other observed engagement metrics. These results suggest that, at least with regards to commenting behavior, Democrats are not engaging with Republican-aligned videos more than Republicans are engaging with Democratic-aligned videos: we find no asymmetric homophily that would explain the ideological skew.
@@
We formalize these analyses in a logistic regression framework (Figure 2H). We conduct a regression at the bot-video recommendation level where the outcome variable indicates whether a video is an ideological mismatch as a function of that bot’s conditioned partisanship, controlling for that bot’s state, the week in which the experiment was conducted, as well as the engagement metrics of the videos and channels watched during the recommendation stage, in addition to those watched in the conditioning stage. We isolate the independent variables with a Variance Inflation Factor (VIF) value less than five (Supplementary Table 11 lists the VIF scores for each of the independent variables). Considering that all conditioning video- and channel-level engagement metrics had a VIF score greater than five, we compute a combined engagement score for those videos using PCA. In Figure 2H, the top-most point shows the logistic regression coefficient on bot conditioning party. While opposition-comment proportion meaningfully predicts a video being mismatched, compared to Republican-conditioned bots, Democratic-conditioned bots remain 2.8 times as likely to receive ideologically mismatched videos even when controlling for opposition comment proportion. We show the full regression table for this model in Supplementary Table 13. Additionally, Supplementary Tables 16-19 include descriptive statistics of the engagement metrics of videos and channels seen by bots during the conditioning and recommendation stages.
Sensitivity Analysis
While we have shown evidence that observable engagement metrics and viewership patterns cannot explain the ideological skew in Republican-aligned content on TikTok, we cannot rule out the possibility that there are unobserved or latent engagement metrics or viewership patterns, known to the TikTok algorithm but unknown to us, that would explain the gap we observe. Instead, we conduct a sensitivity analysis to calculate how strong such an unobserved factor would have to be to fully explain TikTok’s apparent ideological skew.
Across our 14 observed metrics, there are four where Republican-aligned videos outperform Democratic-aligned videos: comments, likes, plays, and shares. If we consider only the optimal linear combination of those four in our weighted sampling approach, they explain 5.7% of the observed ideological gap. Therefore, for any latent or unobserved metric to fully explain that gap, Republican-aligned videos would need to outperform Democratic-aligned videos by 9.4 times as strongly as they do on the aggregation of comments, likes, plays, and shares (see Table 11 in Methods for the required scaling needed on the various engagement metrics analyzed). We argue that such a metric is unlikely to exist. We formalize the assumptions behind this in the Sensitivity Analysis component of the Methods section, in which we experiment with sampling from a variety of distributions.
@@
Indeed, of the top Republican and Democratic channels, the highest four mismatch proportions were of Republican channels (daterightstuff, foxnews, realdonaldtrump, and teamtrump). In contrast, four channels were never shown to a bot of an opposite ideological alignment during our experiments (real.benshapiro, rbreich, jeffjacksonnc, and timwalz). Overall, Republican channels had a significantly higher mismatch proportion than Democratic channels (Chi-squared test; $\chi^{2}=36.5$, $p<0.001$). Figure 3B shows that the official channels of the Republican candidates, Donald Trump and JD Vance, had a significantly higher mismatch proportion than those of Democratic candidates Kamala Harris and Tim Walz (Chi-squared test; $\chi^{2}=28.1$, $p<0.001$). To test whether it is specifically mismatches within top accounts that contribute to the biases observed overall, we compute a logistic regression similar to that shown in Figure 2H, while including another binary channel-level metric designating whether a channel is amongst one of the top 20 channels by follower count. The results of this analysis are in Figure 3C. Here, we again find that the partisan discrepancy observed is most strongly associated with the bot’s partisan conditioning, as well as the opposing comment proportion of a given video. In contrast, we find that top channels were less likely to be recommended as a mismatch, despite the discrepancies seen in Figure 3A and  3B, suggesting that less popular channels were even more likely to be ideologically mismatched. We again considered variables with a VIF value less than 5, as listed in Supplementary Table 12. See Supplementary Table 14 for the logistic regression table containing these results.
@@
Considering this gap, we recompute the ideological skew measured in Figure 2G, once while only considering positive-partisan videos (Pro Democratic and Pro Republican), and again while only considering negative-partisan videos. For each of these sets of videos, we also recompute the expected ideological skew based on the different engagement metrics these videos receive (Supplementary Tables 6 and 7 detail the ideological skews as measured by each of these models for positive-partisan and negative-partisan videos, respectively). The results of these analyses are presented in Figures 4C and  4D for positive- and negative-partisan videos, respectively. Broadly speaking, the Republican skew observed in positive-partisan videos is markedly smaller than that in negative-partisan videos. Formalizing this analysis using a logistic regression (Figure 4E) with video mismatch as an outcome variable suggests that it is specifically Anti Democrat videos which were most likely to be recommended as an ideological mismatch relative to other types of videos. See Supplementary Table 15 for the logistic regression table containing these results.
Survey on TikTok algorithm changes
Figure 5: Survey responses. Regression coefficients of Republican respondents relative to Democratic respondents for different survey questions regarding content on their TikTok feeds. The top four rows are with regards to responses to text-entry questions, while the bottom six rows are with regards to responses to structured scale-based questions. Statistically significant coefficients are highlighted in navy.
While the experimental results above reveal systematic partisan biases in TikTok’s recommendation algorithm, it remains an open question whether these changes are perceptible to real users. Do TikTok users themselves notice shifts in political content on the platform, and do perceptions of these changes vary along partisan lines? To address this question, we conducted a pre-registered survey of 1,000 U.S.-based TikTok users to examine whether they had observed changes in their TikTok feeds over the past year. Our goal was to assess the public’s awareness of algorithmic shifts and determine whether those shifts were interpreted differently based on political affiliation.
Participants were first asked a series of three open-ended, text-entry questions about perceived changes to their TikTok feed. The first question asked generally whether they had noticed any changes in content over the past year; the second focused specifically on changes to political content; and the third asked whether political content had become more positive or more negative. Following these open-ended questions, participants responded to six structured items using a 0–10 scale, measuring their perception of whether their feed had shifted toward more political or less political content, more Democratic or more Republican content, whether it had become more positive or negative, more optimistic or pessimistic, more Pro or Anti Trump, and whether they were seeing more content they agreed or disagreed with politically (see the Survey section of the Methods for more details).
Figure 5 presents the regression coefficients of Republican respondents relative to Democratic respondents across all questions. Specifically, each of the coefficients presented is taken from a separate regression, each of which controls for participant demographics (see Supplementary Table 20 and 21 for the full set of regression results and Supplementary Table 22 for a breakdown of participant demographics). In the free response portion, Republican respondents were less likely than Democrats to mention political content when asked about overall feed changes, and no significant partisan differences were observed when respondents were asked whether they noticed a shift toward Republican or Democratic content in their feed overall. However, when asked about political content specifically, Republicans were significantly more likely to mention increases in co-partisan content and more positive political content.
In the structured response section, there were clear partisan differences as well. Republicans were significantly more likely than Democrats to report that they had seen an increase in party-aligned content, and were more likely to characterize the political content on their feed as positive, optimistic, agreeable, and pro-Trump.
Topic analysis
What do partisan videos discuss, and what topics are most likely to be recommended as an ideological mismatch? To answer these questions, we use methods from [59] to identify the substantive topics covered in a given political video (see details in the Topic Analysis section of the Supplementary Materials). To start, we first isolate topics that were covered in at least 100 unique videos. We then compute the difference between the proportion of Republican- and Democratic-aligned videos on a given topic. Figure 6A illustrates differences in Republican and Democratic coverage of different topics, ordered from most proportionally Republican to most proportionally Democratic. As can be seen, topics stereotypically associated with the Democratic party (e.g., climate change, as well as abortion and reproductive health) emerge as topics with significantly more proportional coverage from Democratic-aligned videos. On the other hand, topics such as immigration, foreign policy, or the Ukraine war were topics more frequently covered by Republican-aligned videos. See Supplementary Table 10 for additional details regarding the distribution of topics within Democratic- and Republican-aligned videos, as well as chi-squared tests comparing these distributions.